there is a triangle that sides represent, 3/y^2, 1/x^2, and 5/x^2. find an expression that represents its perimeter.
Answer by venugopalramana(3286)

there is a triangle that sides represent, 3/y^2, 1/x^2, and 5/x^2. find an expression that represents its perimeter.
LET PERIMETER =P
P=3/Y^2 +1/X^2 +5/X^2 = (3X^2+Y^2+5Y^2)/(X^2*Y^2)=(3(X^2+2Y^2)/(X^2*Y^2)